OpenStudy (anonymous):

Tickets to a play cost $5 at the door and $4 in advance. The theater club wants to raise at least $400 from the play. Write and graph an inequality for the number of tickets the theater club needs to sell. If the club sells 40 tickets in advance, how many do they need to sell at the door to reach their goal?

OpenStudy (anonymous):

let d is the number of customers who will buy tickets at door and a be the number of customers who will book the tickets in advance. total cost is d*5 + a*4 = 400 This is the target. Given a = 40 ==> 5d = 240 => a = 48

OpenStudy (anonymous):

40*4=160, 400-160=240, 240/5=48. 48 tickets at the door to reach the goal.

OpenStudy (anonymous):

how would you graph it?

OpenStudy (anonymous):

its just a straight line with x axis as ticket booking at door (d), y axis as ticket booking in advance (a). 5x+4y = 400.